# Concurrent Fermionic Simulation Gate

### v1

@@ Acknowledgements.
Appendix A Bogoliubov Approach
The Bogoliubov Transformation approach is not so accurate when anharmonicity is non-negligible. Nevertheless, it provides a clear physical picture and intuition. It can be used as a guidance for developing more accurate methods.
Here we use one Bogoliubov transformation to diagonalize the harmonic part of the Hamiltonian. We also assume anharmonicities of the qubits remain unchanged.
The transformation is defined by a unitary transformation $u_{mj}$:
$a_{m}=\sum_{j}u_{mj}\alpha_{j}$
After the transformation the undriven Hamiltonian is diagonalized:
$\begin{split}\tilde{H}&=U^{\dagger}HU\\ &=\sum_{m}\tilde{\omega}_{m}\alpha^{\dagger}_{m}\alpha_{m}+\frac{\delta_{m}}{2}\alpha^{\dagger}_{m}\alpha_{m}(\alpha^{\dagger}_{m}\alpha_{m}-1)\\ \tilde{H}_{d}&=\sum_{mjk}\Omega_{m}u_{mj}^{*}u_{mk}\sin(\nu_{m}t+\phi^{d}_{m})\alpha^{\dagger}_{j}\alpha_{k}\end{split}$ (41)
$\begin{split}\tilde{H}=&\sum_{m}(\tilde{\omega}_{m}+\sum_{j}|u_{jm}|^{2}\Omega_{j}\sin(\omega_{j}^{d}t+\phi_{j}))\alpha^{\dagger}_{m}\alpha_{m}\\ &+\frac{\delta_{m}}{2}\alpha^{\dagger}_{m}\alpha_{m}(\alpha^{\dagger}_{m}\alpha_{m}-1)\\ &+\sum_{mjk}\Omega_{m}\sin(\nu_{m}t+\phi^{d}_{m})u_{mj}^{*}u_{mk}\alpha_{j}^{\dagger}\alpha_{k}\end{split}$
Going to the rotating frame defined by $U_{r}=e^{-i\int\sum_{m}(\tilde{\omega}_{m}+\sum_{j}|u_{jm}|^{2}\Omega_{j}\sin(\nu_{j}t+\phi_{j}))\alpha^{\dagger}_{m}\alpha_{m}dt}$.
The resulting Hamiltonian takes the following form:
$\begin{split}\tilde{H}_{R}=&\sum_{jk}\tilde{g}_{jk}(t)\alpha^{\dagger}_{j}\alpha_{k}\\ &+\sum_{m}\frac{\delta_{m}}{2}\alpha^{\dagger}_{m}\alpha_{m}(\alpha^{\dagger}_{m}\alpha_{m}-1)\end{split}$ (42)
where
$\begin{split}&\tilde{g}_{jk}=\sum_{n_{0},n_{1},n_{2}=-\infty}^{+\infty}\sum_{m=0}^{2}A_{n_{0}n_{1}n_{2}}^{jk}(m)\Omega_{m}u_{mj}^{*}u_{mk}i^{n_{0}+n_{1}+n_{2}}e^{i\phi(t)}\\ &\phi(t)=(\tilde{\omega}_{j}-\tilde{\omega}_{k}-n_{0}\nu_{0}-n_{1}\nu_{1}-n_{2}\nu_{2})t-n_{0}\varphi_{0}-n_{1}\varphi_{1}-n_{2}\varphi_{2}\end{split}$ (43)
$\begin{split}A_{n_{0}n_{1}n_{2}}^{jk}(0)=&\frac{J_{n_{0}+1}(-s_{0}^{jk})+J_{n_{0}-1}(-s_{0}^{jk})}{2}J_{n_{1}}(-s_{1}^{jk})J_{n_{2}}(-s_{2}^{jk})\\ A_{n_{0}n_{1}n_{2}}^{jk}(1)=&J_{n_{0}}(-s_{0}^{jk})\frac{J_{n_{1}+1}(-s_{1}^{jk})+J_{n_{1}-1}(-s_{1}^{jk})}{2}J_{n_{2}}(-s_{2}^{jk})\\ A_{n_{0}n_{1}n_{2}}^{jk}(2)=&J_{n_{0}}(-s_{0}^{jk})J_{n_{1}}(-s_{1}^{jk})\frac{J_{n_{2}+1}(-s_{2}^{jk})+J_{n_{2}-1}(-s_{2}^{jk})}{2}\end{split}$ (44)
$s_{m}^{jk}=\frac{|u_{mj}|^{2}-|u_{mk}|^{2}}{\nu_{m}}\Omega_{m}$
$J_{n}$ is the nth Bessel function of the first kind.
Appendix B Analytical coupling strength by exact diagonalization
@@
$\begin{split}\tilde{H}_{d}&=U^{{\dagger}}H_{d}U\\ &=\sum_{i}\Omega_{i}\sin(\omega_{i}^{d}t+\phi_{i})U^{{\dagger}}a^{\dagger}_{i}a_{i}U\ \end{split}$ (45)
@@
$\begin{split}U^{{\dagger}}a^{\dagger}_{m}a_{m}U=&\sum_{Q_{1}Q_{2}R}N^{m}_{Q_{1}Q_{2}R}|Q_{1}Q_{2}R\rangle\langle Q_{1}Q_{2}R|\\ &+\sum_{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime}\neq Q_{1}Q_{2}R}C^{m}_{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R}|Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime}\rangle\langle Q_{1}Q_{2}R|\\ \end{split}$ (47)
@@
The final Hamiltonian in the rotating frame can be calculated similarly as in the Bogoliubov case.
@@
$\begin{split}G(t)_{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R}&=\sum_{n_{0},n_{1},n_{2}=-\infty}^{+\infty}\sum_{m=0}^{2}A_{n_{0}n_{1}n_{2}}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R}(m)\Omega_{m}C^{m}_{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R}i^{n_{0}+n_{1}+n_{2}=}\\ &*e^{i((\tilde{\omega}_{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime}}-\tilde{\omega}_{Q_{1}Q_{2}R}-n_{0}\nu_{0}-n_{1}\nu_{1}-n_{2}\nu_{2})t-n_{0}\varphi_{0}-n_{1}\varphi_{1}-n_{2}\varphi_{2})}\end{split}$ (52)
$\begin{split}A_{n_{0}n_{1}n_{2}}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R}(0)=&\frac{J_{n_{0}+1}(-s_{0}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R})+J_{n_{0}-1}(-s_{0}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R})}{2}\\ &*J_{n_{1}}(-s_{1}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R})J_{n_{2}}(-s_{2}^{(ijk,npq)})\\ A_{n_{0}n_{1}n_{2}}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R}(1)=&J_{n_{0}}(-s_{0}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R})\\ &*\frac{J_{n_{1}+1}(-s_{1}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R})+J_{n_{1}-1}(-s_{1}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R})}{2}\\ &*J_{n_{2}}(-s_{2}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R})\\ A_{n_{0}n_{1}n_{2}}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R}(2)=&J_{n_{0}}(-s_{0}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R})J_{n_{1}}(-s_{1}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R})\\ &*\frac{J_{n_{2}+1}(-s_{2}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R})+J_{n_{2}-1}(-s_{2}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R})}{2}\end{split}$ (53)
@@
Note that here the definition of coupling strength $G(t)$ is different from $\tilde{g}(t)$ defined in Eq.(42) by the factor $\sqrt{(n_{j}+1)n_{k}}$. This factor is absorbed into the creation and annihilation operators in the Bogoliubov Transformation case.
@@
$\tilde{\omega}_{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime}}-\tilde{\omega}_{Q_{1}Q_{2}R}-n_{0}\nu_{0}-n_{1}\nu_{1}-n_{2}\nu_{2}=0$ (54)
Appendix C Details of crosstalk
Plots of crosstalk effects in Fig.(12).
@@ Appendix C Details of crosstalk
Figure 12: Crosstalk effect of the second drive on the first drive. 12(a) Analytical calculation of static $\bar{g}_{010,100}$ vs $\Omega_{2}$ and $\nu_{2}$ while the first drive is fixed at $\Omega_{1}=150$ MHz, $\nu_{1}=\Delta_{010,100}$. The analytical plot confirms the numerical simulation in Fig.(5). 12(b) Analytical calculation of P10 using the static coupling strength $\bar{g}_{010,100}$. 12(c) Analytical calculation of static $\bar{g}_{010,100}$ vs $\Omega_{2}$ and $\frac{\Omega_{2}}{\nu_{2}}$ while the first drive is fixed at $\Omega_{1}=150$ MHz, $\nu_{1}=\Delta_{010,100}$. Zeros of $J_{0}$ show up at $\frac{\Omega_{2}}{\omega_{2}^{d}}=2.40,5.52,8.65...$. 12(d) Analytical calculation of P100 according to $\bar{g}_{010,100}$ vs $\Omega_{2}$ and $\frac{\Omega_{2}}{\nu_{2}}$.
Appendix D Minimizing Leakage
Plots of 2D scan of optimal drive amplitude in FIig.(13)
(a)Refer to caption(b)
Refer to caption(c)Refer to caption(d)
Figure 13: Optimal $\Omega_{2}$ and $g_{110,020}$ vs $\Omega_{1}$ and $\nu_{2}$ and the corresponding leakage and $P_{110}$. 13(a) Analytically calculated optimal $\Omega_{2}$ vs $\nu_{2}$ and $\Omega_{1}$. The solvable area ($|\nu_{2}-\Delta_{110,020}|<\frac{2\pi}{t_{g}}$) is bounded by the two dark vertical lines indicating $|\nu_{2}-\Delta_{110,020}|=\frac{2\pi}{t_{g}}$. 13(b) Analytically calculated optimal effective coupling $\bar{g}_{110,020}$ vs $\nu_{2}$ and $\Omega_{1}$. 13(c) Leakage vs $\Omega_{1}$ and $\nu_{2}$ from numerical simulation. The amplitude of the second drive $\Omega_{2}$ is determined using the analytical formula. Within the solvable area, the leakage is consistently lower than $1\%$. 13(d) $P_{110}$ vs $\Omega_{1}$ and $\nu_{2}$ from numerical simulation.
Appendix E Pulse Shaping
FIG.14 shows iswap and cphase parameters with pulse shaping.
(a)Refer to caption(b)
(c)Refer to caption(d)
Figure 14: fSim gate with flat-top Gaussian pulse-shaping.  14(a) Fidelity with pulse-shaping. The overall fidelity is improved by almost one order of magnitude. 14(b) Optimization factor for pulse-shaping. The discontinuity in the optimization factor is the reason for the jumps/discontinuity in fidelity, $\theta$ and $\varphi$. 14(c) Numerically calculated $P_{100}$ vs $\Omega_{1}$ and $\nu_{2}$ 14(d) Numerically calculated conditional phase vs $\Omega_{1}$ and $\nu_{2}$. The plot is more irregular due to the optimization needed for pulse-shaping.

### v2

@@ Acknowledgements.
Appendix A Rotating Frame of the Toy Model
The time-dependent Hamiltonian $H(t)$ is given by:
$H_{0}=\begin{pmatrix}\Omega_{2}\sin(\nu_{2}t)&g&0&0&0\\ g&\Delta+\Omega_{1}\sin(\nu_{1}t)&0&0&0\\ 0&0&\Delta+\Omega_{1}\sin(\nu_{1}t)+\Omega_{2}\sin(\nu_{2}t)&\sqrt{2}g&\sqrt{2}g\\ 0&0&\sqrt{2}g&\delta+2\Omega_{2}\sin(\nu_{2}t)&0\\ 0&0&\sqrt{2}g&0&2\Delta+\delta+2\Omega_{1}\sin(\nu_{1}t)\end{pmatrix}$ (41)
[…]
$U=\begin{pmatrix}e^{i\frac{\Omega_{2}}{\nu_{2}}\cos(\nu_{2}t)}&0&0&0&0\\ 0&e^{i[\frac{\Omega_{1}}{\nu_{1}}\cos(\nu_{1}t)-\Delta t]}&0&0&0\\ 0&0&e^{i[\frac{\Omega_{1}}{\nu_{1}}\cos(\nu_{1}t)+\frac{\Omega_{2}}{\nu_{2}}\cos(\nu_{2}t)-\Delta t]}&0&0\\ 0&0&0&e^{i[\frac{2\Omega_{2}}{\nu_{2}}\cos(\nu_{2}t)-\delta t]}&0\\ 0&0&0&0&e^{i[\frac{2\Omega_{1}}{\nu_{1}}\cos(\nu_{1}t)-2\Delta t-\delta t]}\end{pmatrix}$ (42)
The transformed Hamiltonian $H_{I}$ can be easily calculated:
$H_{I}=g\begin{pmatrix}0&e^{i[\frac{\Omega_{1}}{\nu_{1}}\cos(\nu_{1}t)-\frac{\Omega_{2}}{\nu_{2}}\cos(\nu_{2}t)-\Delta t]}&0&0&0\\ h.c.&0&0&0&0\\ 0&0&0&\sqrt{2}e^{i[\frac{\Omega_{2}}{\nu_{2}}\cos(\nu_{2}t)-\frac{\Omega_{1}}{\nu_{1}}\cos(\nu_{1}t)+\Delta t-\delta t]}&\sqrt{2}e^{i[\frac{\Omega_{1}}{\nu_{1}}\cos(\nu_{1}t)-\frac{\Omega_{2}}{\nu_{2}}\cos(\nu_{2}t)-\Delta t-\delta t]}\\ 0&0&h.c.&0&0\\ 0&0&h.c.&0&0\end{pmatrix}$ (43)
where we have used $h.c.$ for Hermitian conjugate.
Simplifying Eq.(43) using the Jacobi–Anger identity (i.e. $\exp({i\alpha\cos\theta})=\sum_{n}i^{n}J_{n}(\alpha)\exp({in\theta})$, one gets the result Eq.(6) in main text.
Appendix B Perturbative treatment of the toy model
Without loss of generality, let us assume that the frequency of Q1 is greater than that of Q2, i.e. $\omega_{1}>\omega_{2}$. For this setup, we can parametrically drive the frequency of Q1 by supplying the required energy to achieve the transition $|01\rangle\leftrightarrow|10\rangle$, i.e. $\nu_{1}=\Delta_{0}$. The second qubit Q2 is parametrically driven to come very close to the resonance between $|11\rangle\leftrightarrow|02\rangle$ transition, i.e. $\nu_{2}=\Delta_{0}+\delta+\varepsilon$ with small $|\varepsilon|$, i.e. $|\varepsilon|\ll\min\{|\Delta_{0}+\delta|,|\Delta_{0}-\delta|\}$. By this assumption, only the slow-varying terms proportional to $\exp(i\varepsilon t)$ contribute to the state evolution in the long run $t\gg 1/(\Delta_{0}\pm\delta)$. The Hamiltonian (6) contains interaction within and outside the Hilbert space’s computational subset. Within the computational subset there is exchange interaction $|10\rangle\langle 01|$ with the strength $gJ_{1}(\Omega_{1}/\Delta_{0})J_{0}(\Omega_{2}/(\Delta_{0}+\delta+\varepsilon))$. Outside of the computational subset the energy level $|11\rangle$ interacts with $|02\rangle$ and $|20\rangle$ by the following coupling strength $\sqrt{2}gJ_{0}(\Omega_{1}/\Delta_{0})J_{1}(\Omega_{2}/(\Delta_{0}+\delta+\varepsilon))e^{i\varepsilon t}$ and $\sqrt{2}gJ_{-2}(\Omega_{1}/\Delta_{0})J_{1}(\Omega_{2}/(\Delta_{0}+\delta+\varepsilon))e^{i\varepsilon t}$, respectively.
One of the aspects of perturbative validity of our result is that the qubits are weakly driven, therefore we consider that the drive amplitude is weak, i.e. ${\Omega_{i}}<\nu_{i}$ for $i=1,2$. Within this limit which makes the arguments of the Bessel function, small, one can show that $J_{-2}J_{1}\ll J_{1}J_{0}$; therefore, the coupling strength of $|11\rangle\langle 20|$ transition turns out to be negligible compared to the other two interaction strengths and can be approximately ignored. Therefore in the interaction picture after rotating-wave-approximation, the BPD Hamiltonian can be simplified as
$H$ $=$ $gJ_{1}(x_{1})J_{0}(x_{2})|10\rangle\langle 01|$ (44)
Appendix C Analytical coupling strength by exact diagonalization
@@
$\begin{split}\tilde{H}_{d}&=U^{{\dagger}}H_{d}U\\ &=\sum_{i}\Omega_{i}\sin(\nu_{i}t+\phi_{i}^{d})U^{{\dagger}}a^{\dagger}_{i}a_{i}U\ \end{split}$ (45)
@@
$U^{{\dagger}}a^{\dagger}_{m}a_{m}U=\sum_{Q_{1}Q_{2}R}N^{m}_{Q_{1}Q_{2}R}|Q_{1}Q_{2}R\rangle\langle Q_{1}Q_{2}R|+\sum_{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime}\neq Q_{1}Q_{2}R}C^{m}_{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R}|Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime}\rangle\langle Q_{1}Q_{2}R|$ (47)
@@
The final Hamiltonian in the rotating frame can be written as:
@@
$\begin{split}G(t)_{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R}&=\sum_{n_{r},n_{1},n_{2}=-\infty}^{+\infty}\sum_{m=r,1,2}A_{n_{r}n_{1}n_{2}}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R}(m)\Omega_{m}C^{m}_{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R}i^{n_{r}+n_{1}+n_{2}}\\ &*e^{i((\tilde{\omega}_{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime}}-\tilde{\omega}_{Q_{1}Q_{2}R}-n_{r}\nu_{r}-n_{1}\nu_{1}-n_{2}\nu_{2})t-n_{r}\varphi_{r}-n_{1}\varphi_{1}-n_{2}\varphi_{2})}\end{split}$ (52)
$\begin{split}A_{n_{r}n_{1}n_{2}}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R}(r)=&\frac{J_{n_{r}+1}(-s_{r}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R})+J_{n_{r}-1}(-s_{r}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R})}{2}\\ &*J_{n_{1}}(-s_{1}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R})J_{n_{2}}(-s_{2}^{(ijk,npq)})\\ A_{n_{r}n_{1}n_{2}}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R}(1)=&J_{n_{r}}(-s_{r}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R})\\ &*\frac{J_{n_{1}+1}(-s_{1}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R})+J_{n_{1}-1}(-s_{1}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R})}{2}\\ &*J_{n_{2}}(-s_{2}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R})\\ A_{n_{r}n_{1}n_{2}}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R}(2)=&J_{n_{r}}(-s_{0}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R})J_{n_{1}}(-s_{1}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R})\\ &*\frac{J_{n_{2}+1}(-s_{2}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R})+J_{n_{2}-1}(-s_{2}^{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime},Q_{1}Q_{2}R})}{2}\end{split}$ (53)
@@
Note that here the definition of coupling strength $G(t)$ is different from usual definition of $g$ by the factor $\sqrt{(n_{j}+1)n_{k}}$.
@@
$\tilde{\omega}_{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime}}-\tilde{\omega}_{Q_{1}Q_{2}R}-n_{r}\nu_{r}-n_{1}\nu_{1}-n_{2}\nu_{2}=0$ (54)
Appendix D Non-RWA effects on iSWAP angles and conditional phases
The iSWAP angles and conditional phases are not so much affected by non-RWA effects as fidelity. We show the simulation results in FIG. (13).
@@ Appendix D Non-RWA effects on iSWAP angles and conditional phases
Figure 13: iSWAP and CPHASE patterns with non-RWA terms included. 13(a) iSWAP pattern without pulse-shaping. The iSWAP pattern is shifted towards larger $\Omega_{1}$. This is likely because the effective coupling between $|100\rangle$ and $|010\rangle$ is reduced by the non-RWA terms. 13(b) CPHASE pattern without pulse-shaping. The CPHASE pattern remains mostly unchanged. 13(c) iSWAP pattern with pulse-shaping. 13(d) CPHASE pattern with pulse-shaping. Similar as the patterns without pulse-shaping.
Appendix E fSim gates via DC mixed with AC drive.
The simultaneous transition scheme is very flexible. Instead of using two parametric drives for two transitions, one can activate one transition via a DC drive ( two levels tuned on resonance) and drive the other transition via a parametric drive. Here we use the two-qutrit toy model to demonstrate the idea.
We drive the iSWAP transition by tuning $|01\rangle$ and $|10\rangle$ into resonance and drive the CPHASE transition via a parametric drive. For simplicity, let’s assume we can directly control the coupling strength $g(t)=g_{0}+\Omega\sin(\omega t)$. The time-dependent coupling strength between two qutrits has a DC part $g_{0}$ and a parametrically driven part $\Omega\sin(\omega t)$. Similar as Eq.(4), we can then write down the Hamiltonian in the Hilbert space spanned by $\{|01\rangle,|10\rangle,|11\rangle,|02\rangle,|20\rangle\}$:
$H(t)=\begin{pmatrix}0&g(t)&0&0&0\\ g(t)&0&0&0&0\\ 0&0&0&\sqrt{2}g(t)&\sqrt{2}g(t)\\ 0&0&\sqrt{2}g(t)&\delta_{2}&0\\ 0&0&\sqrt{2}g(t)&0&\delta_{1}\end{pmatrix}$ (55)
We numerically solve the time-depedent Schrödinger equation under this Hamiltonian. Similar behaviours as BPD dynamics can be observed. Given a fixed gate time, there are ”leakage centers” corresponding to population of $|11\rangle$ being coherently driven out of the computational space. Away from those centers, leakage is small. The conditional phase is tunable by varying drive amplitude and frequency. iSWAP rate is mainly controlled by the DC drive $g_{0}$. In the simulation, we have chosen gate time $t_{g}$ and DC drive $g_{0}$ so that iSWAP angle $\theta$ is around $90^{\circ}$, see Fig. (14)
Figure 14: Numerical simulation of DC-AC driven fSim gates. We assume two qutrits have same anharmonicity $\delta_{1}/2\pi=\delta_{2}/2\pi=-200$ MHz. Gate time $t_{g}$ is chosen to be around 25 ns. The DC drive $g_{0}/2\pi$ is $10$ MHz. 14(a)Population of $|11\rangle$. The two leakage centers occur when AC drive frequency $\omega$ is matching with anharmonicity causing resonant oscillations among $|11\rangle$, $|02\rangle$ and $|20\rangle$. 14(b) Conditinal phase of $|11\rangle$. A $\pi$ phase can be realized under resonant AC drive, corresponding to a CZ operation. The zero conditional phase trajectories are marked with red dashed lines. 14(c) Population of $|11\rangle$. Similar as the BPD case, there is some crosstalk from the AC drive to the iSWAP rate. This can be understood as baseband modulation.